you give a salesperson $20 for a purchase and receive change. Use your function to find the amount of change from a $4.50 purchase. what is the function rule
Answer by josmiceli(19441)

c = change from $20
p = amount of purchase
c = 20 - p
c = 20 -(4.50)
c = 15.50
c = 20 - p is the correct way to show the function because c depends on p
and not the other way around. p is the INdependant variable
c is the DEpendant variable. All that means is- you've got to know p
before you know c.